a boy is 10 years old and his mother is 37 years old. In how many years time will the mother be twice as old as her son?
Found 2 solutions by Alan3354, Edwin McCravy:
Answer by Alan3354(69443)

a boy is 10 years old and his mother is 37 years old. In how many years time will the mother be twice as old as her son?
--------------
The difference will always be 27.
--> when they're 27 & 54.

Answer by Edwin McCravy(20055)

a boy is 10 years old and his mother is 37 years old. In how many years time
will the mother be twice as old as her son?
Two ways to do it.  With and without algebra.

Without algebra.

She will be twice as old as he is when his age is the difference between their
ages, which is 37-10 or 27.  He will be 27 in 27-10 or 17 more years.

Answer: 17 years.  Checking: He'll be 27 and she'll be 54.

----------------------------

With algebra:

In x years he will be 10+x
In x years his mom will be 37+x

His mom will then equal twice his age.

37+x will equal 2 times 10+x.

37+x = 2(10+x)

37+x = 20 + 2x
  17 = x

Answer: 17 years. Checking: He'll be 27 and she'll be 54.

Edwin